PLZ HELP!  The number of chips of different colors in Vicky's bag is shown below:


5 blue chips

11 pink chips

9 white chips

Vicky takes out a chip from the bag randomly without looking. She replaces the chip and then takes out another chip from the bag. What is the probability that Vicky takes out a blue chip in both draws?

A5/25 X 5/25 = 25/625
B5/25 + 5/25 = 10/25
C5/25 X 4/24 = 20/600
D5/25 + 4/24 = 220/600